OpenStudy (anonymous):

Use an inequality to solve the problem. Be sure to show the inequality and all of your work. The color guard is making new triangular flags that must have a base of 18 inches to fit on their flagpoles. What is the maximum height of the triangular flags, if they want to use a maximum of 180 square inches of cloth? Is this right: 1/2(18)(h)-180-a^2

OpenStudy (anonymous):

\[\frac{1}{2}(18\times h)\leq 180\] \[\therefore h\leq 20\]

OpenStudy (anonymous):

As the area of the flag must be less than or equal to 180 square inches.

OpenStudy (anonymous):

I understand better now. I'm about to try and solve. brb with the results.

OpenStudy (anonymous):

Familiarize: We were given the base of the flag pole at 18 in and the area sum of 180 square in. Translate: Since the sum of the triangle is less then or equal to 180, we have 1/2(18h) ≤ 180 Carry out: Solve Equation 1/2(18×h) ≤ 180 18h * 1/2 ≤ 180 37h/2 *2/37 ≤ 180 * 2/37 h ≤ 20